Given log b 3= 0.6826 and log b 7 + 1.2091, evaluate log b 63.
Answer by mukhopadhyay(490)

log b 63 = log(baseb)[3*3*7]=log(baseb)3 + log(baseb)3 +log(baseb)7
=2(0.6826) + 1.2091 = 2.5743